At a baseball game, a vender sold a combined total of 230 sodas and hot dogs. The number of hot dogs sold was 38 less than the number of sodas sold. Find the number of sodas sold and the number of hot dogs sold.
I am having trouble setting up the problem. Could you please help me?

Answer by mananth(16946)

x= Number of hamburgers
y= Number of sodas

1 x + 1 y = 230 .............1
Total value
1 x -1 y = 38 .............2
Eliminate y
Add the two equations
2 x = 268
/ 2
x = 134
plug value of x in (1)
1 x + 1 y = 230
134 + y = 230
y = 230 -134
y = 96
y = 96
x= 134 Number of hamburgers
y= 96 Number of sodas
m.ananth@hotmail.ca